OpenStudy (anonymous):

The graph below models the path of a golf ball after it was hit. Write an equation in vertex form that represents the path of the ball.

OpenStudy (anonymous):

y=-3/50(x-50)^2+150*I think. y=-3/2(x-150)^2+50 y=-3/20(x-100)^2+150 y=-50/30(x-50)^2+150